# Critical temperatures of two dimensional magnets beyond linear spin wave theory: application to CrI3, MPS3 (M=Ni, Mn, Fe) and CrSBr

### v1

@@ II.2.3 Magnetization and Critical temperature
In the Green’s function approach, the (sublattice) magnetization is given by [14, 9, 15]
@@
Since $\omega_{\eta,\mathbf{q}}$ is proportional to $\langle S^{{\prime\prime}z}\rangle$, the expression effectively no longer depends on it. This is a generalization of the critical temperature of a Bravais lattice derived in our previous work [9] to non-Bravais lattices with equivalent spin sites. For the RPA+CD approach, the critical temperature must be obtained from a numerical (self-consistent) calculation of the magnetization curve.
@@ II.3 Critical temperatures in the limit of small anisotropies
$\omega^{\mathrm{FM}}_{\mathbf{q}}=\langle S^{{\prime\prime}z}\rangle(bq^{2}+\Delta_{\mathrm{FM}})$ (62)
@@
Note that there is an additional term proportional to $(cq^{2}+\Delta_{\mathrm{AFM}}^{2})^{-1/2}$ originating from the eigenvectors when evaluating Eq. (54) for anti-ferromagnets.
Figure 1: The magnon dispersion of monolayer CrI3 calculated using bulk INS parameters from Ref. [19]. A DM interaction of 0.09 meV in the out-of-plane direction opens a topological gap at the high symmetry point K. The inset shows the magnetic lattice of CrI3.
@@ II.4 Dipole interactions
Since the dipolar interaction is inherently long ranged, the dipole tensor in Eq. (72) must be evaluated to sufficiently large distances until the summation converges. Unlike the single-ion anisotropy, the dipolar anisotropy is dispersive ($q$-dependent) and therefore manifests itself differently in the magnon spectrum.
@@ III.1 CrI3
CrI3 was the first material shown to exhibit long-range ferromagnetic order in the monolayer limit [1]. The Cr atoms form a honeycomb lattice and exhibits strong easy-axis uniaxial anisotropy. In fact, the three-fold rotational symmetry associated with each magnetic site renders the single-ion anisotropy tensor diagonal with $\mathrm{K}_{a}^{xx}=\mathrm{K}_{a}^{yy}$, and we are free to set these in-plane components to zero. The single-ion anisotropy is thus completely characterized by $\mathrm{K}_{a}^{zz}$. Since its discovery, there has been considerable interest in this material - theoretical as well as experimental [20, 21]. The Cr atoms are in the 3+ oxidation state and are well represented by $S=3/2$ sites in a Heisenberg model. Since CrI3 is ferromagnetic, there is no reduction in the saturation magnetization, which is simply given by $S^{(0)}_{a}=3/2$ as well. The Heisenberg exchange and anisotropy parameters extracted from inelastic neutron scattering (INS) experiments of bulk CrI3 [19] are listed in Tab. 1. As we are interested in the monolayer limit, the interlayer exchange is not included in our calculations, but we assume that the Heisenberg parameters for a single monolayer are well approximated by those of the bulk material. The Dzyaloshinskii–Moriya (DM) interaction in CrI3 is absent in the nearest neighbor bond (due to an inversion center), but appears between next nearest neighbors. It is reported to be relatively weak compared to the exchange constants, with a magnitude of $0.09$ meV [19], but gives rise to a sizable gap between the acoustic and optical magnon branches. At the level of linear spin-wave theory it is only the out-plane component that opens such a gap and the remaining components are typically disregarded although they are not forbidden by symmetry in CrI3. The gap between the magnon branches is topological (finite Chern number) and has been discussed extensively in the literature [19, 22, 23]. In the present tensor notation, these out-of-plane DM component enter as intra-sublattice components $\mathrm{J}^{xy}_{aaij}/2=-\mathrm{J}^{yx}_{aaij}/2$, where $i$ and $j$ denote neighboring cells. Since single-ion anisotropy contributes differently to the magnon dispersion for HP and RPA treatments, the reported experimental values depend on which method the dispersion was fitted to. We have thus recalculated the uniaxial single-ion anisotropy constant, $\mathrm{K}^{zz}_{a}$, directly from the experimentally observed spin-wave gap, obtaining a value of 0.184 meV for RPA+CD (and HP) and 0.123 meV for RPA. This recalculation ensures a consistent treatment of anisotropy within our formalism and ensures that the calculated spinwave dispersion agrees with the experimental one regardless of the methodology. We note that a similar analysis was performed in our previous work on CrI3 [15], but in that case we employed exchange constants extracted from optical magnon energies and the spin-wave gap of bilayer CrI3 reported in Ref. [24]. In contrast, here we base our analysis on the full magnon spectrum of bulk CrI3 measured via INS experiments, providing a more comprehensive and accurate parameter set.
@@
The magnon dispersion along the high-symmetry path $\Gamma$–K–M (with INS parameters from Ref. [19]) is shown in Fig. 1 and the corresponding critical temperatures obtained using MC, HP, RPA, and RPA+CD methods are summarized in Tab. 3. It is observed that the RPA+CD approach yields the closest agreement to the experimental critical temperature, which is consistent with previous work [15] based on parameters from bilayer CrI3.
@@
Figure 2: The magnon dispersion relations of monolayer MPS3 (M=Ni, Fe, Mn) anti-ferromagnets with Heisenberg parameters taken from INS [25]. The inset shows the magnetic lattices with the corresponding ground state magnetic order.
@@ III.2 MPS3
The MPS3 (M=Ni, Mn, Fe) compounds comprise a family of layered transition-metal thiophosphates that crystallize with a honeycomb arrangement of transition-metal ions within each layer. Here, we consider NiPS3, FePS3 and MnPS3, which gives rise to layered anti-ferromagnetic ordering with uniaxial anisotropy [26, 27, 28, 29, 30]. NiPS3 and FePS3 exhibit a zig-zag magnetic ordering in the $ab$-plane with an in-plane ordering vector $\mathbf{Q}=[1/2,0]$, where spins in the unit cell are parallel with $S=1$ and $S=2$ respectively. MnPS3, on the other hand exhibits Néel order, where all the nearest neighbors are anti-parallel and is represented by a $\mathbf{Q}=[0,0]$ in-plane ordering vector with anti-parallel spins in the primitive unit cell and $S=5/2$. The exchange and anisotropy parameters of these compounds have been studied extensively in the literature using both experimental [28, 30, 25, 29] and first principles approaches [31, 32]. Similar to the case of CrI3, symmetry dictates that the single-ion anisotropy is completely characterized by $\mathrm{K}_{a}^{zz}=\mathrm{K}^{zz}$. For the purpose of benchmarking our methods, we use the Heisenberg parameters obtained from the bulk INS data reported in Ref. [25], and the values are tabulated in Tab. 1. The calculated magnon dispersions for the three compounds are shown in fig. 2, along with the magnetic structure. Since these materials are anti-ferromagnets, the exchange constants ($J_{n}$) from the experimental fit performed with linear spin wave theory differ from RPA/RPA+CD by a factor of $S_{a}/\langle S^{{\prime\prime}z}_{a}\rangle$. To avoid confusion, we have retained the original unscaled exchange constants and assumed $S^{(0)}_{a}=S_{a}$ in our calculations.
@@
Table 1: In-plane exchange and anisotropy constants of bulk CrI3 [19] and MPS3 (M=Ni, Fe, Mn) [25] taken from INS experiments ($\mathrm{K}^{xx}=\mathrm{K}^{yy}=0$). All numbers are meV. The anisotropy constants are calculated from the spin-wave dispersion by fitting to either RPA or HP/RPA+CD (shown in brackets).
@@ III.3 CrSBr
CrSBr is a layered van der Waals bonded anti-ferromagnetic semiconductor, in which the individual ferromagnetic layers are coupled anti-ferromagnetically [34, 35, 36]. The ferromagnetic monolayer has been exfoliated from bulk and exhibits a remarkably high Curie temperature of 146 K [37, 38, 39, 40]. The Cr3+ magnetic ions form a rectangular lattice with a ground state spin of $S=3/2$. In addition, it is notable for exhibiting triaxial magnetic anisotropy, with distinct easy, intermediate, and hard axes [41, 42].
@@
Table 2: In-plane exchange constants (in meV) of bulk CrSBr taken from Ref. [43]. $J_{6}$ is omitted as it corresponds to interlayer exchange in the layered compound. The anisotropy parameters (in meV) are taken from spectroscopic measurements [41] and are converted (using isotropic invariance) such that the hard axis anisotropy component ($\mathrm{K}_{a}^{zz}$) is zero. The values are listed for RPA/RPA+CD/HP, which are obtained by fitting the predicted dispersion relations to the experimental one.
The Heisenberg parameters of CrSBr used in this work are compiled from two experimental references. The isotropic exchange interactions of bulk CrSBr are taken from Ref. [43] where the anisotropy was reported too weak to be measured. Since anisotropy is a crucial parameter for magnetic order, we use the values reported in Ref. [41] where the hard and intermediate anisotropy parameters are obtained from anti-ferromagnetic resonance experiments. In CrSBr, the hard axis is along the $c$-axis, the intermediate is the $a$-axis and the easy axis is the $b$-axis. The point group of the monolayer is mmmmmm and the mirror symmetries enforce all off-diagonal elements of the single-ion anisotropy tensors to vanish. We will again use the isotropic invariance of Eq. (2) to set the hard axis components $\mathrm{K}_{a}^{zz}$ to zero and we convert the values stated in Ref. [41] to our conventions with a full characterization given by the components $\mathrm{K}_{a}^{xx}$ and $\mathrm{K}_{a}^{yy}$.
@@
CrSBr 56 95 151 150 146
Table 3: Calculated critical temperatures (in K) of the materials considered in this work. The experimental values marked with a star are critical temperatures in the bulk compounds, while the remaining ones (including all calculations) are for monolayers.
@@ IV Conclusion
We have extended the theoretical approaches for calculating critical temperatures of single-$Q$ 2D magnetic materials - emphasizing the role of single-ion anisotropy. We studied the HP approach, and observed that normal ordering of the fourth order bosonic operators yield second-order corrections to the zero temperature magnon energies. These corrections capture the correct behavior of uniaxial single-ion anisotropy for ferromagnets, but not for anti-ferromagnets or for general systems with triaxial anisotropy. The corrections also fail to conserve the isotropic invariance of single-ion anisotropy terms because of an incomplete bosonic expansion of the transverse spin components. We then studied the RPA decoupling scheme applied to the single ion anisotropy terms and it was shown that RPA conserves the isotropic invariance but fails to reproduce the correct spin-wave gap - in particular the irrelevance of single-ion anisotropy for $S=1/2$. Finally, we extended Callen’s decoupling scheme to include the single-ion anisotropy tensor of any single-$Q$ system. In contrast to RPA, CD captures the correct spin-wave gap and irrelevance of single-ion anisotropy for $S=1/2$ ferromagnets. Moreover, in contrast to HP it conserves the isotropic invariance and its accuracy is not restricted to uniaxial cases. On the basis of fundamental theoretical arguments, we expect CD to provide the best description of single-ion anisotropy. For magnon dispersion relations, we thus adopt a hybrid RPA+CD approach, using Callen’s decoupling for the single-ion anisotropy and RPA for the remaining interactions. This combination has also been shown to yield the best agreement with quantum Monte Carlo results for ferromagnets [16].
@@
In summary, we presented a comprehensive analysis of the role of single-ion anisotropy in calculating thermally excited magnon energies and critical temperatures in two dimensional magnets. Among the methods considered, the RPA+CD approach showed the best agreement with experiments, establishing it as the best method for predicting critical temperatures in low-dimensional systems. This approach comprises a simple and versatile method that is easily applicable and is expected to significantly improve the accuracy of either classical Monte Carlo simulations or linear spin wave theory, which seems to be the most common approaches in the literature. The framework can also be used to analyze the influence of complex spin interactions, such as anisotropic exchange, Dzyaloshinskii-Moriya interaction, dipole-dipole interactions and higher order terms like biquadratic exchange. As such, we believe that the RPA+CD approach provides the most reliable route for understanding and predicting thermal properties of 2D materials.
@@ Appendix A Exchange contribution to the HP Hamiltonian
Note that the expressions here are slightly different from the ones presented in Ref. [9] because we have separated the single-ion anisotropy tensor from $\mathrm{J}_{ab\mathbf{q}}$. Here the $q$-space exchange interactions are defined as
@@
where $n_{\mathrm{B}}(\omega_{n,\mathbf{q}})$ is the Bose distribution and $\mathrm{T}_{\mathbf{q}}$ is the canonical matrix that transforms $\mathbf{a}_{\mathbf{q}}$ to a basis that (para)diagonalizes the Hamiltonian (20) [13, 9]
@@
The procedure is described in more detail in Ref. [9].

### v2

@@ II.2.3 Magnetization and Critical temperature
In the Green’s function approach, the (sublattice) magnetization is given by [Callen1963, Helimag3d2025, BeyondRPA2025]
@@
Since $\omega_{\eta,\mathbf{q}}$ is proportional to $\langle S^{{\prime\prime}z}\rangle$, the expression effectively no longer depends on it. This is a generalization of the critical temperature of a Bravais lattice derived in our previous work [Helimag3d2025] to non-Bravais lattices with equivalent spin sites. For the RPA+CD approach, the critical temperature must be obtained from a numerical (self-consistent) calculation of the magnetization curve.
@@ II.3 Critical temperatures in the limit of small anisotropies
$\omega^{\mathrm{FM}}_{\mathbf{q}}=\langle S^{{\prime\prime}z}\rangle(bq^{2}+\Delta_{\mathrm{FM}}).$ (63)
@@
Note that for antiferomagnets an additional term proportional to $(cq^{2}+\Delta_{\mathrm{AFM}}^{2})^{-1/2}$ arises from the eigenvectors $U_{a\eta\mathbf{q}}U^{-1}_{\eta a\mathbf{q}}$ when the magnon number (55) is evaluated.
Figure 1: The magnon dispersion of monolayer CrI3 calculated using bulk INS parameters from Ref. [Chen-2021-cri3]. A DM interaction of 0.09 meV in the out-of-plane direction opens a topological gap at the high symmetry point K. The inset shows the magnetic lattice of CrI3.
@@ II.4 Dipole interactions
Since the dipolar interaction is inherently long ranged, the dipole tensor in Eq. (73) must be evaluated to sufficiently large distances until the summation converges. Unlike the single-ion anisotropy, the dipolar anisotropy is dispersive ($q$-dependent) and therefore manifests itself differently in the magnon spectrum. The long range nature of the interaction also implies that the Mermin-Wagner theorem does not apply, and easy-plane order in 2D can be stabilized at finite temperatures by the dipole-dipole interaction even if it does not open a gap in the magnon spectrum [bruno_spin-wave_1991].
@@ III.1 CrI3
CrI3 was the first material shown to exhibit long-range ferromagnetic order in the monolayer limit [Huang2017]. The Cr atoms form a honeycomb lattice and exhibits strong easy-axis uniaxial anisotropy. In fact, the three-fold rotational symmetry associated with each magnetic site renders the single-ion anisotropy tensor diagonal with $\mathrm{K}_{a}^{xx}=\mathrm{K}_{a}^{yy}$, and we are free to set these in-plane components to zero. The single-ion anisotropy is thus completely characterized by $\mathrm{K}_{a}^{zz}$. Since its discovery, there has been considerable interest in this material - theoretical as well as experimental [Soriano2020, Chen2018a]. The Cr atoms are in the 3+ oxidation state and are well represented by $S=3/2$ sites in a Heisenberg model. Since CrI3 is ferromagnetic, there is no reduction in the saturation magnetization, which is simply given by $S^{(0)}_{a}=3/2$ as well. The Heisenberg exchange and anisotropy parameters extracted from inelastic neutron scattering (INS) experiments of bulk CrI3 [Chen-2021-cri3] are listed in Tab. 1. As we are interested in the monolayer limit, the interlayer exchange is not included in our calculations, but we assume that the Heisenberg parameters for a single monolayer are well approximated by those of the bulk material. The Dzyaloshinskii–Moriya (DM) interaction in CrI3 is absent in the nearest neighbor bond (due to an inversion center), but appears between next nearest neighbors. It is reported to be relatively weak compared to the exchange constants, with a magnitude of $0.09$ meV [Chen-2021-cri3], but gives rise to a sizable gap between the acoustic and optical magnon branches. At the level of linear spin-wave theory it is only the out-plane component that opens such a gap and the remaining components are typically disregarded although they are not forbidden by symmetry in CrI3. The gap between the magnon branches is topological (finite Chern number) and has been discussed extensively in the literature [Chen-2021-cri3, Mook2021, Yaroslav2016]. In the present tensor notation, these out-of-plane DM component enter as intra-sublattice components $\mathrm{J}^{xy}_{aaij}/2=-\mathrm{J}^{yx}_{aaij}/2$, where $i$ and $j$ denote neighboring cells. Since single-ion anisotropy contributes differently to the magnon dispersion for HP and RPA treatments, the reported experimental values depend on which method the dispersion was fitted to. We have thus recalculated the uniaxial single-ion anisotropy constant, $\mathrm{K}^{zz}_{a}$, directly from the experimentally observed spin-wave gap, obtaining a value of 0.184 meV for RPA+CD (and HP) and 0.123 meV for RPA. This recalculation ensures a consistent treatment of anisotropy within our formalism and ensures that the calculated spinwave dispersion agrees with the experimental one regardless of the methodology. We note that a similar analysis was performed in our previous work on CrI3 [BeyondRPA2025], but in that case we employed exchange constants extracted from optical magnon energies and the spin-wave gap of bilayer CrI3 reported in Ref. [Cenker2021]. In contrast, here we base our analysis on the full magnon spectrum of bulk CrI3 measured via INS experiments, providing a more comprehensive and accurate parameter set.
@@
The magnon dispersion along the high-symmetry path $\Gamma$–K–M (with INS parameters from Ref. [Chen-2021-cri3]) is shown in Fig. 1 and the corresponding critical temperatures obtained using MC, HP, RPA, and RPA+CD methods are summarized in Tab. 3. It is observed that the RPA+CD approach yields the closest agreement to the experimental critical temperature, which is consistent with previous work [BeyondRPA2025] based on parameters from bilayer CrI3.
@@
Figure 2: The magnon dispersion relations of monolayer MPS3 (M=Ni, Fe, Mn) anti-ferromagnets with Heisenberg parameters taken from INS [Wildes2018]. The inset shows the magnetic lattices with the corresponding ground state magnetic order.
@@ III.2 MPS3
The MPS3 (M=Ni, Mn, Fe) compounds comprise a family of layered transition-metal thiophosphates that crystallize with a honeycomb arrangement of transition-metal ions within each layer. Here, we consider NiPS3, FePS3 and MnPS3, which gives rise to layered anti-ferromagnetic ordering with uniaxial anisotropy [Vasudevan1992, Jernberg1984, Wildes1998, Wildes2015, Wildes2016]. NiPS3 and FePS3 exhibit a zig-zag magnetic ordering in the $ab$-plane with an in-plane ordering vector $\mathbf{Q}=[1/2,0]$, where spins in the unit cell are parallel with $S=1$ and $S=2$ respectively. MnPS3, on the other hand exhibits Néel order, where all the nearest neighbors are anti-parallel and is represented by a $\mathbf{Q}=[0,0]$ in-plane ordering vector with anti-parallel spins in the primitive unit cell and $S=5/2$. The exchange and anisotropy parameters of these compounds have been studied extensively in the literature using both experimental [Wildes1998, Wildes2016, Wildes2018, Wildes2015] and first principles approaches [Olsen_2021, Kim2021-mps3]. Similar to the case of CrI3, symmetry dictates that the single-ion anisotropy is completely characterized by $\mathrm{K}_{a}^{zz}=\mathrm{K}^{zz}$. For the purpose of benchmarking our methods, we use the Heisenberg parameters obtained from the bulk INS data reported in Ref. [Wildes2018], and the values are tabulated in Tab. 1. The calculated magnon dispersions for the three compounds are shown in fig. 2, along with the magnetic structure. Since these materials are anti-ferromagnets, the exchange constants ($J_{n}$) from the experimental fit performed with linear spin wave theory differ from RPA/RPA+CD by a factor of $S_{a}/\langle S^{{\prime\prime}z}_{a}\rangle$. To avoid confusion, we have retained the original unscaled exchange constants and assumed $S^{(0)}_{a}=S_{a}$ in our calculations.
@@
Table 1: In-plane exchange and anisotropy constants of bulk CrI3 [Chen-2021-cri3] and MPS3 (M=Ni, Fe, Mn) [Wildes2018] taken from INS experiments ($\mathrm{K}^{xx}=\mathrm{K}^{yy}=0$). All numbers are meV. The anisotropy constants are calculated from the spin-wave dispersion by fitting to either RPA or HP/RPA+CD (shown in brackets).
@@ III.3 CrSBr
CrSBr is a layered van der Waals bonded anti-ferromagnetic semiconductor, in which the individual ferromagnetic layers are coupled anti-ferromagnetically [Katscher1966-crsbr, Beck1990-crsbr, Goser1990-crsbr]. The ferromagnetic monolayer has been exfoliated from bulk and exhibits a remarkably high Curie temperature of 146 K [Lee2021-CrSBr, Telfor2020-crsbr, Ziebel2024-CrSBr, Lopez2022-crsbr]. The Cr3+ magnetic ions form a rectangular lattice with a ground state spin of $S=3/2$. In addition, it is notable for exhibiting triaxial magnetic anisotropy, with distinct easy, intermediate, and hard axes [Bae2022-CrSBr-K, Yang2021-crsbr-dft].
@@
Table 2: In-plane exchange constants (in meV) of bulk CrSBr taken from Ref. [Scheie2022]. $J_{6}$ is omitted as it corresponds to interlayer exchange in the layered compound. The anisotropy parameters (in meV) are taken from spectroscopic measurements [Bae2022-CrSBr-K] and are converted (using isotropic invariance) such that the hard axis anisotropy component ($\mathrm{K}_{a}^{zz}$) is zero. The values are listed for RPA/RPA+CD/HP, which are obtained by fitting the predicted dispersion relations to the experimental one.
The Heisenberg parameters of CrSBr used in this work are compiled from two experimental references. The isotropic exchange interactions of bulk CrSBr are taken from Ref. [Scheie2022] where the anisotropy was reported too weak to be measured. Since anisotropy is a crucial parameter for magnetic order, we use the values reported in Ref. [Bae2022-CrSBr-K] where the hard and intermediate anisotropy parameters are obtained from anti-ferromagnetic resonance experiments. In CrSBr, the hard axis is along the $c$-axis, the intermediate is the $a$-axis and the easy axis is the $b$-axis. The point group of the monolayer is mmmmmm and the mirror symmetries enforce all off-diagonal elements of the single-ion anisotropy tensors to vanish. We will again use the isotropic invariance of Eq. (2) to set the hard axis components $\mathrm{K}_{a}^{zz}$ to zero and we convert the values stated in Ref. [Bae2022-CrSBr-K] to our conventions with a full characterization given by the components $\mathrm{K}_{a}^{xx}$ and $\mathrm{K}_{a}^{yy}$.
@@
CrSBr 56 91 151 150 146
Table 3: Calculated critical temperatures (in K) of the materials considered in this work. The experimental values marked with a star are critical temperatures in the bulk compounds, while the remaining ones (including all calculations) are for monolayers. The MC simulations are performed with a $30\times 30$ super cell. The HP and Green’s function calculations are performed self-consistently by iteratively converging the magnetization at each temperature step.
@@ IV Conclusion
We have extended the theoretical approaches for calculating critical temperatures of single-$Q$ 2D magnetic materials - emphasizing the role of single-ion anisotropy. We studied the HP approach, and observed that normal ordering of the fourth order bosonic operators yield second-order corrections to the zero temperature magnon energies. These corrections capture the correct behavior of uniaxial single-ion anisotropy for ferromagnets, but not for anti-ferromagnets or for general systems with triaxial anisotropy. We then studied the RPA decoupling scheme applied to the single ion anisotropy terms and it was shown that RPA fails to reproduce the correct spin-wave gap - in particular the irrelevance of single-ion anisotropy for $S=1/2$. Finally, we extended Callen’s decoupling scheme to include the single-ion anisotropy tensor of any single-$Q$ system. In contrast to RPA, CD captures the correct spin-wave gap and irrelevance of single-ion anisotropy for $S=1/2$ ferromagnets. Moreover, in contrast to HP, its accuracy is not restricted to uniaxial cases. On the basis of fundamental theoretical arguments, we expect CD to provide the best description of single-ion anisotropy. For magnon dispersion relations, we thus adopt a hybrid RPA+CD approach, using Callen’s decoupling for the single-ion anisotropy and RPA for the remaining interactions. This combination has also been shown to yield the best agreement with quantum Monte Carlo results for ferromagnets [Frobich2006].
@@
In summary, we presented a comprehensive analysis of the role of single-ion anisotropy in calculating thermally excited magnon energies and critical temperatures in two dimensional magnets. Among the methods considered, the RPA+CD approach showed the best agreement with experiments, establishing it as the best method for predicting critical temperatures in low-dimensional systems. This approach comprises a simple and versatile method that is easily applicable and is expected to significantly improve the accuracy of either classical Monte Carlo simulations or linear spin wave theory, which seems to be the most common approaches in the literature. Although no explicit benchmarking was performed for non-collinear states, all of these methods are directly applicable to such systems, provided they retain the single-$Q$ symmetry. The framework can also be used to analyze the influence of complex spin interactions, such as anisotropic exchange, Dzyaloshinskii-Moriya interaction, dipole-dipole interactions and higher order terms like biquadratic exchange. As such, we believe that the RPA+CD approach provides the most reliable route for understanding and predicting thermal properties of 2D materials.
@@ Appendix A Exchange contribution to the HP Hamiltonian
The linear and cubic order bosonic terms do not contribute to the dispersion as they can be removed by a linear transformation. Note that the expressions here are slightly different from the ones presented in Ref. [Helimag3d2025] because we have separated the single-ion anisotropy tensor from $\mathrm{J}_{ab\mathbf{q}}$. Here the $q$-space exchange interactions are defined as
@@
where $n_{\mathrm{B}}(\omega_{n,\mathbf{q}})$ is the Bose distribution and $\mathrm{T}_{\mathbf{q}}$ is the canonical matrix that transforms $\mathbf{a}_{\mathbf{q}}$ to a basis that (para)diagonalizes the Hamiltonian (20) [Toth2015, Helimag3d2025]
@@
The procedure is described in more detail in Ref. [Helimag3d2025].